OpenStudy (anonymous):

How many solutions does Cos^-1(0) have? My answer is 1...is this correct?

OpenStudy (ash2326):

Yeah you're right

OpenStudy (anonymous):

\[ \frac \pi2 \] if you mean the one valued function \[ \cos^{-1}(0) \]

OpenStudy (anonymous):

1 is not right.

OpenStudy (anonymous):

@eliassaab , is that how many solutions Cos^-1(0) can have? Or something else?

OpenStudy (ash2326):

If \[y = \cos^{-1} 0\] then there is one solution for y but if we have \[ \cos y=0\] and we need to find the solution of this then there are innumerable solutions

OpenStudy (anonymous):

Yeah it's how many solutions there are for the entire notation Cos^-1(0)...so it's infinite?

OpenStudy (ash2326):

@missfitz5172 you're correct. It has only 1 solution because \[ \cos^{-1} x\] is defined like that its range is [0, \(\pi\)] so only one solution exist at \(\pi/2\)

OpenStudy (anonymous):

@ash2326 oh ok that makes sense! thank you very much!

OpenStudy (ash2326):

You got it??

OpenStudy (anonymous):

yeah! (:

OpenStudy (ash2326):

Welcome to Open Study:) Glad to help you:D